OpenStudy (anonymous):

Seven subtracted from 3 times a number is equal to 9 more than 5 times the number. What is the number?

OpenStudy (anonymous):

3n-7=5n+9

OpenStudy (anonymous):

yes now what @zimmah

OpenStudy (anonymous):

solve for n

OpenStudy (anonymous):

for example, subtract 9 from both the left and the right hand side to start with

OpenStudy (anonymous):

3n-7=5n+9 \(3n-7\color{red}{-9} = 5n+9\color{red}{-9}\) makes 3n-16 = 5n

OpenStudy (anonymous):

then subtract 3n from both sides

OpenStudy (anonymous):

guess you should be able to work it out from there

OpenStudy (anonymous):

now what @zimmah

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (anonymous):

sorry i was busy with something else in the maintime

OpenStudy (anonymous):

here are some tools that might help you http://nlvm.usu.edu/en/nav/frames_asid_201_g_4_t_2.html https://www.youtube.com/watch?v=k6pq60KnWEo